# First-Order Sparse Convex Optimization: Better Rates with Sparse Updates

### v2

@@ 4 Some Numerical Evidence
We set the dimension $n=3000$, the radius of the $\ell_{1}$ ball to $R=10$. The optimal solution ${\mathbf{x}}^{*}$ is drawn at random so that each non-zero entry (the non-zero locations are also randomized) is set to ±$\pm\frac{R}{\textrm{nnz}({\mathbf{x}}^{*})}$ with equal probability, where $\textrm{nnz}({\mathbf{x}}^{*})$ is set in advanced, and for each level of sparsity we run 10 i.i.d. experiments (where in each one we resample ${\mathbf{x}}^{*}$). Accordingly, the figure below presents the average approximation error over 10 i.i.d. experiments.
When implementing our Algorithm 1, we assumed the exact sparsity of ${\mathbf{x}}^{*}$ is known, i.e., we set $s=\textrm{nnz}({\mathbf{x}}^{*})$. Additionally, observing that the theoretical choice of step-size corresponding to Theorem 1, $\eta=\frac{\alpha_{2}}{6\beta_{1}{}s}$ (see Section 3.1) is too conservative in practice, we used a larger step-size $\tilde{\eta}=\frac{\alpha_{2}}{\beta_{1}{}s}$, when computing the point ${\mathbf{v}}_{t}$ on each iteration $t$. Additionally, instead of using again the step-size $\tilde{\eta}$ when taking the convex combination to produce the next iterate ${\mathbf{x}}_{t+1}$, we took a slightly larger step-size $2\tilde{\eta}$ (which is still in (0,1)(0,1)( 0 , 1 )). For the V-FISTA method we observed that even trying to slightly tune the prescribed theoretical step-size makes the method divergent, and so we stuck to the original implementation in [2].
From Figure 1, we can clearly see that Algorithm 1 is significantly faster for highly-sparse instances, strongly outperforming even a state-of-the-art accelerated method, while as ${\mathbf{x}}^{*}$ becomes less-sparse, this advantage vanishes.
(a) $\textrm{nnz}({\mathbf{x}}^{*})=10$
(b) $\textrm{nnz}({\mathbf{x}}^{*})=30$
(c) $\textrm{nnz}({\mathbf{x}}^{*})=50$
Figure 1: Comparison of Algorithm 1 and the V-FISTA algorithm for Problem (1).
@@ 5 Open Questions
We conclude with two open questions which we believe are of interest. First, it is interesting if our approach could be extended to also handle norm-regulairzed problems instead of norm-constrained problems. Second, it is interesting if our approach could be accelerated, i.e., to yield worst-case number of iterations which scales only with $\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}$ while maintaining the sparse updates.
@@ Acknowledgments
Appendix A Proof of Lemma 1
@@
Denote $k=\textrm{nnz}({\mathbf{x}})$. Let $S\subseteq[n]$ be the set of indices for non-zero entries in ${\mathbf{x}}$. Let ${\mathbf{x}}_{1},{\mathbf{y}}_{1}\in\mathbb{R}^{k}$ be the restrictions of ${\mathbf{x}}$ and ${\mathbf{y}}$ to the entries indexed by $S$, respectively. Accordingly, let ${\mathbf{x}}_{2},{\mathbf{y}}_{2}\in\mathbb{R}^{n-k}$ be the restrictions of ${\mathbf{x}}$ and ${\mathbf{y}}$ to the entries not indexed by $S$. It holds that,
@@
Appendix B Proof of Lemma 2

### v3

@@ 4 Numerical Evidence
We present simple numerical simulations on the $\ell_{1}$-constrained Problem (1). Similarly to the example in Eq. (3), we set $f({\mathbf{x}})=\frac{1}{2}({\mathbf{x}}-{\mathbf{x}}^{*})^{\top}({\mathbf{I}}+3\cdot\mathbf{1}\mathbf{1}^{\top})({\mathbf{x}}-{\mathbf{x}}^{*})$, where $\mathbf{1}$ denotes the all-ones vector. For this problem we have that the $\ell_{2}$-quadratic growth (in fact, strong convexity) constant is $\alpha_{2}=1$, and the $\ell_{2}$ and $\ell_{1}$ Lipschitz continuity constants of the gradient are $\beta_{2}=1+3n$ and $\beta_{1}=4$, respectively. In all experiments we set the dimension to $n=1000$, the radius of the $\ell_{1}$ ball to $R=10$. The optimal solution $\mathbf{x}$$^{*}$ is drawn at random so that each non-zero entry (the non-zero locations are also randomized) is set to $\pm\frac{R}{\mathrm{nnz}({\mathbf{x}}^{*})}$ with equal probability, where $\mathrm{nnz}({\mathbf{x}}^{*})$ is set in advance, and for each level of sparsity we run 10 i.i.d. experiments (where in each one we resample $\mathbf{x}$$^{*}$) and display the average approximation error vs. number of iterations in Figure 1.
Aside from our Algorithm 1, we implemented three benchmarks: I. the away-steps Frank-Wolfe algorithm with line-search that for this setting has worst-case convergence rate $O\left({\frac{\beta_{1}sR^{2}}{\alpha_{2}}\log{1/\epsilon}}\right)$ (see Section A), II. a state-of-the-art Euclidean accelerated gradient method for smooth and strongly convex optimization known as V-FISTA [2], which has worst-case convergence rate $O((1-\sqrt{\alpha_{2}/\beta_{2}})^{t})$ (see Theorem 10.42 in [2]), and III. a restarted non-Euclidean accelerated gradient method with hard-thresholding steps as described in Section B which has convergence rate $\tilde{O}\left({\sqrt{\frac{\beta_{1}{}s\ln(n)}{\alpha_{2}}}\log{1/\epsilon}}\right)$ (per Remark 4 we implemented Nesterov’s Method 5.6 [23] as our non-Euclidean accelerated gradient method). For the V-FISTA method we observed that even trying to slightly tune the prescribed theoretical step-size makes the method divergent, and so we stuck to the original implementation in [2]. For the accelerated non-Euclidean method we set the length of each epoch (the parameter $K$ in Algorithm 3) to $K=\left\lceil{25\sqrt{\frac{\beta_{1}{}s}{\alpha_{2}}}}\right\rceil$ which gave the best results.
As for our Algorithm 1, we implemented three variants, all of them use exact line-search to determine the parameter $\gamma_{t}$, but with different choice of step-size $\eta$: I. using the theoretical step-size $\eta$ in Theorem 1 which, as discussed in Section 3.1 amounts to setting $\eta=\frac{\alpha_{2}}{48\beta_{1}s}$, II. using a larger manually-tuned heuristic step-size $\eta=\frac{\alpha_{2}}{2\beta_{1}s}$, and III. using an automatically-tuned step-size, as discussed in Remark 1, by trying on each iteration all step-sizes $\eta=\frac{2^{i}\alpha_{2}}{48\beta_{1}s}$ for $i\in\{0,1,\dots,5\}$ and taking the one which results in the largest decrease in function value. Note that for $i=0$ we get the theoretical step-size and thus this method is guaranteed to converge with at least the rate established in Theorem 1.
For all methods that require the knowledge of $s$ (an upper-bound on $\mathrm{nnz}({\mathbf{x}}^{*})$), which includes our algorithm and the restarted non-Euclidean accelerated gradient method, we set for simplicity $s=\mathrm{nnz}({\mathbf{x}}^{*})$.
Let us comment on the results displayed in Figure 1. First, we see that our Algorithm 1, when implemented exactly according to theory and with a fixed step-size, does not seem to work well in practice (perhaps with the exception of the case $\mathrm{nnz}({\mathbf{x}}^{*})=10$). This seems to be due to the small theoretical step-size (note the $1/48$ factor in the theoretical choice of $\eta$), which may be an artifact of our analysis. Once we use a manually-tuned step-size or the automatically-tuned step-size (which as stated above has the theoretical guarantee of Theorem 1), we observe state-of-the-art performance, clearly outperforming all benchmarks. Somewhat surprisingly, the restarted non-Euclidean accelerated gradient method fails to substantially outperform the Euclidean V-FISTA method, despite having a theoretically superior convergence rate when $\mathbf{x}$$^{*}$ is sparse. Similarly, except for the highly-sparse setting $\mathrm{nnz}({\mathbf{x}}^{*})=10$, the away-steps Frank-Wolfe also fails to outperform V-FISTA.
(a) $\mathrm{nnz}({\mathbf{x}}^{*})=10$
(b) $\mathrm{nnz}({\mathbf{x}}^{*})=40$
(c) $\mathrm{nnz}({\mathbf{x}}^{*})=100$
Figure 1: Empirical convergence of various algorithms for Problem (1).
@@ 5 Open Questions
We conclude with two open questions which we believe are of interest. First, it is interesting if our approach could be extended to also handle norm-regularized problems instead of norm-constrained problems. Second, it is interesting if our approach could be accelerated, i.e., to yield worst-case number of iterations which scales only with $\sqrt{\frac{\beta_{1}s}{\alpha_{2}}}$ while maintaining the sparse updates.
@@ Acknowledgments
Appendix A Away-Steps Frank-Wolfe for Problem (1)
In this section we show that the away-steps Frank-Wolfe method [14, 20] has a linear convergence rate of the form $O\left({\frac{\beta_{1}sR^{2}}{\alpha_{2}}\log{1/\epsilon}}\right)$ for Problem (1), assuming all optimal solutions lie on the boundary of the $\ell_{1}$ ball.
Theorem 2 (Away-Steps Frank-Wolfe for Problem (1)).
Consider Problem (1) and suppose that for any optimal solution $\mathbf{x}$$^{*}$ it holds that $\mathrm{nnz}({\mathbf{x}}^{*})\leq s$ and $\|{{\mathbf{x}}^{*}}\|_{1}=R$. Then, the Away-steps Frank-Wolfe algorithm (Algorithm 1 in [20], see also [14]) finds an $\epsilon$-approximated solution (in function value) after at most $O\left({\frac{\beta_{1}sR^{2}}{\alpha_{2}}\log\frac{\beta_{1}R^{2}}{\epsilon}}\right)$ iterations. In particular, when $f({\mathbf{x}})=\frac{1}{2}{\mathbf{x}}^{\top}{\mathbf{A}}{\mathbf{x}}+{\mathbf{b}}^{\top}{\mathbf{x}}$ (${\mathbf{A}}\succeq 0$), the overall runtime is $O\left({\frac{ns\beta_{1}R^{2}}{\alpha_{2}}\log\frac{\beta_{1}R^{2}}{\epsilon}}\right)$.
@@ Appendix A Away-Steps Frank-Wolfe for Problem (1)
The theorem follows directly by instantiating Corollary 3.33 in [5], where we take the underlying norm to be the $\ell_{1}$ norm, and by using the fact that, as noted in [5], the $\ell_{1}$-pyramidal width of the $\ell_{1}$ ball is $1$. This leads to an $O\left({\frac{\beta_{1}R^{2}}{\alpha_{1}}\log\frac{\beta_{1}R^{2}}{\epsilon}}\right)$ upper-bound on the number of iterations to reach $\epsilon$-error, where $\alpha_{1}$ is the quadratic growth constant w.r.t. the $\ell_{1}$-norm. The proof follows since the assumption that all optimal solutions are on the boundary of the ball together with Lemma 1, implies that $\alpha_{1}\geq\frac{\alpha_{2}}{4s}$. ∎
Appendix B Accelerated Rates with Dense Updates — Recovering the rates in [15]
In this section we give a concise description of the accelerated linear convergence rates reported in [15], which scale only with $\sqrt{\frac{\beta_{1}{}s}{\alpha_{2}}}$, but which rely on non-sparse updates. While the methods in [15] were developed for stochastic optimization, here we are concerned with a purely deterministic setting. For ease of presentation, throughout this section we will focus on Problem (1) (optimization over $\ell_{1}$ balls in $\mathbb{R}^{n}$) only, however the extension to Problem (2) (optimization over a matrix nuclear norm ball) is straightforward by replacing the entrywise $\ell_{1}$ norm with the matrix nuclear norm. The method is a straightforward combination of a restarted-accelerated non-Euclidean proximal gradient method due to Nesterov [23] with sparse projection steps of the iterates.
Throughout this section we let $\mathcal{B}_{1}(R)$ denote the origin-centered $\ell_{1}$-ball of radius $R$ in $\mathbb{R}^{n}$. For ease of presentation, here (and only here) we assume that Problem (1) admits a unique optimal solution $\mathbf{x}$$^{*}$ (this is also assumed in [15]) and that $\mathrm{nnz}({\mathbf{x}}^{*})\leq s<<n$, however, here we do not require that the optimal solutions lie on the boundary of the $\ell_{1}$ ball.
The following result is well-known, see for instance [15] .
Lemma 3.
Given ${\mathbf{x}}_{0}\in\mathcal{B}_{1}(R)$, consider the prox function
$d({\mathbf{x}}):=\frac{1}{2}e\log{n}\cdot{}n^{(p-1)(2-p)/p}\|{{\mathbf{x}}-{\mathbf{x}}_{0}}\|_{p}^{2}\quad\textrm{for}\quad p=1+1/\log{n}.$ (9)
Then, $d({\mathbf{x}})$ is $1$-strongly convex over $\mathcal{B}_{1}(R)$ w.r.t. the norm $\|{\cdot}\|_{1}$, and for any ${\mathbf{y}}\in\mathcal{B}_{1}(R)$, $d({\mathbf{y}})\leq\frac{e^{2}\log{n}}{2}\|{{\mathbf{y}}-{\mathbf{x}}_{0}}\|_{1}^{2}$.
Algorithm 2 Nesterov’s accelerated gradient method (see Method 3.11 in [23])
1: input: initialization point ${\mathbf{x}}_{0}\in\mathcal{B}_{1}(R)$, $\ell_{1}$-smoothness parameter $\beta_{1}$
2: define the function $d({\mathbf{z}})$ as in Eq. (9) and set $\sigma=1$
3:for $k=0,1\dots$ do
4:  ${\mathbf{y}}_{k}\leftarrow\arg\min_{{\mathbf{y}}\in\mathcal{B}_{1}(R)}\langle{{\mathbf{y}}-{\mathbf{x}}_{k},\nabla{}f({\mathbf{x}}_{k})}\rangle+\frac{\beta_{1}}{2}\|{{\mathbf{y}}-{\mathbf{x}}_{k}}\|_{1}^{2}$
5:  ${\mathbf{z}}_{k}\leftarrow\arg\min_{{\mathbf{z}}\in\mathcal{B}_{1}(R)}\left\{{\frac{\beta_{1}}{\sigma}d({\mathbf{z}})+\sum_{i=0}^{k}\frac{i+1}{2}\langle{{\mathbf{z}}-{\mathbf{x}}_{i},\nabla{}f({\mathbf{x}}_{i})}\rangle}\right\}$
6:  ${\mathbf{x}}_{k+1}\leftarrow\frac{2}{k+3}{\mathbf{z}}_{k}+\frac{k+1}{k+3}{\mathbf{y}}_{k}$
7:end for
Theorem 3 (Theorem 2 in [23]).
When applied to Problem (1), the sequence $({\mathbf{y}}_{k})_{k\geq 0}$ produced by Algorithm 2 satisfies
$\forall k\geq 0:\qquad f({\mathbf{y}}_{k})-f({\mathbf{x}}^{*})\leq\frac{4\beta_{1}d({\mathbf{x}}^{*})}{(k+1)(k+2)}.$
Algorithm 3 Restarted accelerated gradient method
1: input: sparsity parameter $s\in[n]$, epoch length $K$
[…]
3:for $t=0,1\dots$ do
4:  ${\mathbf{x}}_{t+1}\leftarrow$ the point ${\mathbf{y}}_{K}$ computed by Algorithm 2 after $K$ iterations, when initialized with the point $\widehat{{\mathbf{x}}}_{t}=\arg\min_{{\mathbf{x}}\in\mathbb{R}^{n}:~\mathrm{nnz}({\mathbf{x}})\leq s}\|{{\mathbf{x}}-{\mathbf{x}}_{t}}\|_{2}$ {i.e., $\widehat{{\mathbf{x}}}_{t}$ equals ${\mathbf{x}}_{t}$ on the $s$ largest (in absolute value) entries, and zero elsewhere}
5:end for
Theorem 4.
Consider applying Algorithm 3 to Problem (1) with $K\geq\sqrt{\frac{64s\beta_{1}e^{2}\ln(n)}{\alpha_{2}}}-1$. Then,
$\forall t\geq 0:\qquad f({\mathbf{x}}_{t+1})-f({\mathbf{x}}^{*})\leq\frac{\alpha_{2}R^{2}}{s}2^{-(t+3)}.$
Thus, to reach an $\epsilon$-approximated solution (in function value) the number of calls to the first-order oracle of $f$ is $O\left({\sqrt{\frac{\beta_{1}{}s\ln(n)}{\alpha_{2}}}\ln\left({\frac{\alpha_{2}R^{2}}{s\epsilon}}\right)}\right)$.
Proof.
Fix some iteration $t\geq 0$ of Algorithm 3. Let $d_{t}({\mathbf{x}})$ denote the function $d({\mathbf{x}})$, as defined in Eq. (9), with ${\mathbf{x}}_{0}=\widehat{{\mathbf{x}}}_{t}$ — the initialization point in iteration $t$ of Algorithm 3.
$\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t+1}}\|_{1}^{2}$ $\underset{(a)}{\leq}2s\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t+1}}\|_{2}^{2}$
$\leq 4s\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t+1}}\|_{2}^{2}+4s\|{{\mathbf{x}}_{t+1}-\widehat{{\mathbf{x}}}_{t+1}}\|_{2}^{2}$
$\underset{(b)}{\leq}8s\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{t+1}}\|_{2}^{2}\underset{(c)}{\leq}\frac{16s}{\alpha_{2}}\left({f({\mathbf{x}}_{t+1})-f({\mathbf{x}}^{*})}\right)$
$\underset{(d)}{\leq}\frac{64s\beta_{1}d_{t}({\mathbf{x}}^{*})}{\alpha_{2}(K+1)^{2}}\underset{(e)}{\leq}\frac{32s\beta_{1}e^{2}\ln(n)\cdot{}\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{1}^{2}}{\alpha_{2}(K+1)^{2}},$
where (a) holds since $\mathbf{x}$$^{*}$ is $s$-sparse by assumption and $\widehat{{\mathbf{x}}}_{t+1}$ is $s$-sparse by construction, (b) holds by definition of $\widehat{{\mathbf{x}}}_{t+1}$, (c) holds due to the $\ell_{2}$-quadratic growth assumption, (d) follows from Theorem 3 (when applied with the initialization point $\widehat{{\mathbf{x}}}_{t}$ and as a consequence the prox function is $d_{t}(\cdot)$ as defined above), and (e) follows from upper-bounding $d_{t}({\mathbf{x}}^{*})$ using Lemma 3.
Thus, for $K\geq\sqrt{\frac{64s\beta_{1}e^{2}\log(n)}{\alpha_{2}}}-1$, and noting that $\widehat{{\mathbf{x}}}_{0}={\mathbf{x}}_{0}$, we can deduce that
$\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{1}^{2}\leq 2^{-t}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{0}}\|_{1}^{2}.$
Using Theorem 3 and Lemma 3 again, this yields
$f({\mathbf{x}}_{t+1})-f({\mathbf{x}}^{*})$ $\leq\frac{4\beta_{1}d_{t}({\mathbf{x}}^{*})}{(K+1)^{2}}\leq\frac{2\beta_{1}e^{2}\log(n)\|{{\mathbf{x}}^{*}-\widehat{{\mathbf{x}}}_{t}}\|_{1}^{2}}{(K+1)^{2}}$
$\leq\frac{\alpha_{2}}{s}2^{-(t+5)}\|{{\mathbf{x}}^{*}-{\mathbf{x}}_{0}}\|_{1}^{2}\leq\frac{\alpha_{2}R^{2}}{s}2^{-(t+3)}.$
Remark 4.
In the derivations above, for the sake of being concrete, we chose to present Nesterov’s Method 3.11 from [23] as Algorithm 2. A different choice could have been Method 5.6 from the same paper for which Theorem 3 still holds, however, the practical implementation of Method 5.6 might be easier since it requires that both prox computations are w.r.t. the same function $d(\cdot)$, instead of one w.r.t. the $\ell_{1}$ norm and one w.r.t. the function $d(\cdot)$ as in Algorithm 2.
Appendix C Proof of Lemma 1
Proof.
@@ Appendix C Proof of Lemma 1
Denote $k=\mathrm{nnz}({\mathbf{x}})$. Let $S\subseteq[n]$ be the set of indices for non-zero entries in $\mathbf{x}$}. Let ${\mathbf{x}}_{1},{\mathbf{y}}_{1}\in\mathbb{R}^{k}$ be the restrictions of $\mathbf{x}$} and $\mathbf{y}$} to the entries indexed by $S$, respectively. Accordingly, let ${\mathbf{x}}_{2},{\mathbf{y}}_{2}\in\mathbb{R}^{n-k}$ be the restrictions of $\mathbf{x}$} and $\mathbf{y}$} to the entries not indexed by $S$. It holds that,
@@
Appendix D Proof of Lemma 2